# Possible Liquid-Nitrogen-Temperature Superconductivity Driven by Perpendicular Electric Field in the Single-Bilayer Film of La3Ni2O7 at Ambient Pressure

### v2

@@ Abstract
The discovery of high-temperature superconductivity (SC) (HTSC) in pressurized La3Ni2O7 with critical temperature $T_{c}$ higher than the boiling point of liquid nitrogen has aroused a surge in the exploration of HTSC in the Ruddlesden-Popper phase multilayer nickelates. Very recently, SC is found in the La3Ni2O7 ultrathin film grown on the SrLaAlO4 substrate with $T_{c}$ above the McMillan limit ($\approx 40\text{ K}$) at ambient pressure (AP), allowing various experimental investigation on the pairing mechanism in this material. It is now eager to enhance the $T_{c}$ of La3Ni2O7 at AP. Here we propose that an imposed strong perpendicular electric field can strongly enhance the $T_{c}$ in the single-bilayer film of La3Ni2O7 at AP. The physics underlying this proposal is clear and simple. Under strong electric field, the layer with lower potential energy will accept electrons flowing from the other layer to fill in the Ni-$3d_{x^{2}-y^{2}}$ orbitals in this layer, as the nearly half-filled Ni-$3d_{z^{2}}$ orbital in this layer cannot accommodate more electrons. With the enhancement of the filling fraction in the $3d_{x^{2}-y^{2}}$ orbitals in this layer, the interlayer $s$-wave pairing will be subjected to the pair-breaking effect and be suppressed, but the intralayer $d$-wave pairing in this layer is promptly and strongly enhanced, which mimics the cuprates. Our combined simplified one-orbital study and comprehensive two-orbital one under the mean-field treatment and the density matrix renormalization group approach consistently verify this idea and yield that an imposed voltage of about $0.1\sim 0.2$ volt between the two layers is enough to realize HTSC with $T_{c}$ above the boiling point of liquid nitrogen in this single bilayer at AP. Our results appeal for experimental verification.
@@
The discovery of superconductivity (SC) with critical temperature $T_{c}$ above the boiling point of liquid nitrogen ($\approx 77\text{ K}$) in the pressurized La3Ni2O7 [1, 2, 3, 4, 5, 6, 7, 8, 9] has attracted great interests [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144]. This discovery has sparked the exploration of high-temperature SC (HTSC) in Ruddlesden-Popper phase multilayer nickelates, resulting in the discovery of SC in the pressurized La4Ni3O10 [39, 40, 41, 42, 43, 44, 45], which together with the previously synthesized infinite-layer nickelates Nd1-xSrxNiO2 [46, 47, 48, 49] have established a new family of SC other than cuprates and iron-based superconductors. However, the high pressure (HP) circumstance not only strongly hinders the experimental detection of the samples but also brings difficulties in the application of SC in industry. Very recently, the La3Ni2O7 ultrathin film with a few layers of unit cell grown on the SrLaAlO4 (SLAO) substrate has been grown by two different groups independently and SC with $T_{c}$ above the McMillan limit ($\approx$40 K) has been detected at ambient pressure (AP)[132, 133, 134], allowing various experimental investigation of the pairing mechanism in this material, attracting a lot of interests [135, 136, 137, 138, 139, 140, 141, 142, 143, 144]. It is now eager to enhance the $T_{c}$ of this material at AP. Here we propose a viable approach to realize $T_{c}$ above the boiling point of liquid nitrogen in the La3Ni2O7 single-bilayer film at AP.
@@
Presently, the pairing mechanism in the La3Ni2O7, either in the bulk material under HP [74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 104, 105, 106, 107, 108, 125, 126, 127, 128] or in the ultrathin film at AP [135, 138, 140, 143], is still under debate. Density-functional-theory (DFT) based first-principle calculations have suggested that the low-energy orbitals are mainly Ni-$3d_{z^{2}}$ and $3d_{x^{2}-y^{2}}$, which are nearly half- and quarter- filled [50, 51, 52, 53, 54, 55, 56, 57, 58, 61]. Various experiments have revealed the strongly-correlated characteristic of the material [14, 16, 22, 24, 25, 28, 31, 32]. Particularly, the optical study reveals significant reduction of the electron kinetic energy which places the system in the proximity of the Mott phase [25]; the angle-resolved photoemission spectroscopy uncovers strong band renormalization caused by electron correlation [28]; the linearly temperature-dependent resistivity suggests “strange-metal” behavior [2]. Therefore, we take a strong-coupling view of the system. Under the strong Hubbard repulsion, the nearly half-filled $3d_{z^{2}}$ electrons can almost be viewed as localized spins. Therefore, the main carrier of SC should be the $3d_{x^{2}-y^{2}}$ electrons, which subject to the in-plane superexchange interaction just mimic the 50% hole-doped cuprates. However, it is a problem how HTSC can emerge under such a high doping level. The key physics lies in the important role played by the $3d_{z^{2}}$ orbitals. Through strong interlayer superexchange, the $3d_{z^{2}}$ electrons form interlayer pairing. The interlayer superexchange or the interlayer pairing of the $3d_{z^{2}}$ electrons can be transmitted to the $3d_{x^{2}-y^{2}}$ electrons through the Hund’s rule [79, 80, 82, 87, 92, 93, 95, 96, 97, 98, 100, 107, 108] or the nearest-neighbor (NN) hybridization [65, 83, 86, 87, 91, 98, 105, 107, 108] or both. Under such view, the role of pressure in enhancing the $T_{c}$ lies in the enhancement of the interlayer superexchange, the inter-orbital hybridization, or both.
@@
Figure 1: (a) Schematic diagram for the dominant hopping integrals and superexchange interactions between the $E_{g}$ orbitals in La3Ni2O7. (b) Schematic diagram illustrating that the Hund’s rule coupling transmits the interlayer superexchange interaction $J_{\perp}$ between the $3d_{z^{2}}$ orbitals to the effective one $\tilde{J}_{\perp}$ between the $3d_{x^{2}-y^{2}}$ orbitals.
@@
In this work, we propose an alternative approach to realize HTSC with $T_{c}$ above the boiling point of liquid nitrogen in the ultrathin film of La3Ni2O7 at AP. Here we consider the thinnest limit, i.e. a single bilayer film of La3Ni2O7. We can impose a perpendicular electric field, say pointing upward, in this single bilayer, so that electrons from the top layer will flow to the bottom layer. These electrons will fill the $3d_{x^{2}-y^{2}}$ orbitals in the bottom layer as the nearly half-filled $3d_{z^{2}}$ orbitals there cannot accommodate more electrons. The enhancement of the bottom-layer $3d_{x^{2}-y^{2}}$ electron number will first suppress the interlayer $s$-wave SC due to mismatch of the electron numbers between the two layers, and then promptly lead to the intra-bottom-layer $d$-wave SC with strongly enhanced $T_{c}$. To test this idea, we have performed a combined simplified single orbital study and a comprehensive two orbital study, which consistently yield that a voltage of experimentally achievable levels (around $0.1\sim 0.2$ volt predicted by the mean-field calculations) between the two layers is enough to induce $d$-wave SC with $T_{c}$ above the boiling point of liquid nitrogen in the bottom layer. Intriguingly, the $d$-wave SC carried by the bottom layer $3d_{x^{2}-y^{2}}$ electrons coexists with the interlayer $s$-wave pseudo gap carried by the $3d_{z^{2}}$ electrons in the mixing ratio of $1:\mathrm{i}$, breaking time-reversal symmetry. Our proposal potentially provides a viable approach to realize HTSC with $T_{c}$ above the boiling point of liquid nitrogen in the single bilayer film of La3Ni2O7.
@@ I.1 Consideration and a Simplified Study
The La3Ni2O7 ultrathin film grown on the SLAO substrate form a bilayer square lattice [137, 135]. As illustrated in Fig. 1 (a), the leading hopping integrals are the interlayer hopping of the $3d_{z^{2}}$ electrons $t_{\perp}$ and the intralayer NN hopping of the $3d_{x^{2}-y^{2}}$ electrons $t_{\parallel}$. Under strong Hubbard $U$, these hopping terms can induce the effective superexchange $J_{\perp}$ and $J_{\parallel}$ through $J\approx\frac{4t^{2}}{U}$. Under the Hund’s rule coupling $J_{H}$, the spins of the two orbitals are inclined to be parallel aligned, as illustrated in Fig. 1 (b), which partly transmits the interlayer superexchange $J_{\perp}$ between the $3d_{z^{2}}$ orbitals to the $3d_{x^{2}-y^{2}}$ orbitals as $\tilde{J}_{\perp}=\alpha J_{\perp}$ with $\alpha\in(0,1)$. In addition, there exists intralayer NN- hybridization $t_{xz}$ between the two orbitals. As shown in Fig. 2 (a, b), the nearly quarter-filled $3d_{x^{2}-y^{2}}$ electrons subject to $J_{\parallel}$ and $\tilde{J}_{\perp}$ form interlayer-dominant pairing [79].
Figure 2: (a) Particle numbers of the four $E_{g}$ orbitals within an unit cell without electric field. (b) The dominant pairing configuration for (a). (c) Schematic diagram showing how the electrons flow under the perpendicular electric field $\bm{\varepsilon}$ pointing upward. (d) The dominant pairing configuration for (c).
Now let us turn on the upward electric field $\bm{\varepsilon}$, forcing electrons downward, as shown in Fig. 2 (c, d). In the top layer, since the $3d_{z^{2}}$ orbitals host larger density of state (DOS) than the $3d_{x^{2}-y^{2}}$ orbitals, they will donate more electrons. Most of these donated electrons will fill the $3d_{x^{2}-y^{2}}$ orbitals in the bottom layer, as the nearly half-filled $3d_{z^{2}}$ orbitals there cannot accommodate more electrons. A minority of the donated electrons can also be accepted by the top-layer $3d_{x^{2}-y^{2}}$ orbitals.
@@
Figure 3: (a) The pairing amplitude $\tilde{\Delta}$ (in unit of $t_{\parallel}$) as function of the bottom-layer particle number per site $n_{bx}$. Different pairing symmetries are distinguished by color. (b) The $T_{c}$ as function of $n_{bx}$, in comparison with $0.42\tilde{\Delta}$ for the $d$-wave and $0.9\tilde{\Delta}$ for the $s$-wave regime. Inset: the spinon pairing temperature $T_{\mathrm{pair}}$ and the holon condensation temperature $T_{\mathrm{BEC}}$ as function of $n_{bx}$. In (a,b), we set $J_{\parallel}=0.4t_{\parallel}$ and $\tilde{J}_{\perp}=(1-\delta_{tz})\times 1.3J_{\parallel}$. (c)-(d) The pairing configurations of the $s$-wave and $d$-wave, respectively.
Even with doped holes under $\bm{\varepsilon}$, the top-layer $3d_{z^{2}}$ electrons cannot carry SC: Firstly, lacking pairing interaction, they cannot form intralayer pairing. Secondly, although they can pair with the localized bottom-layer $3d_{z^{2}}$ electrons, such pairs cannot coherently move, only resulting in the pseudo-gap. Therefore, the SC under $\bm{\varepsilon}$ can only be carried by the $3d_{x^{2}-y^{2}}$ orbitals. As the filling fractions of the $3d_{x^{2}-y^{2}}$ orbitals in the two layers are different, their Fermi levels are relatively shift, leading to mismatch of their Fermi surfaces (FSs), which will suppress their interlayer pairing. The bottom-layer $3d_{x^{2}-y^{2}}$ orbitals will form $d$-wave SC, mimicsing the cuprates, as shown in Fig. 2 (d). When $\bm{\varepsilon}$ is strong enough so that the filling fraction of the bottom-layer $3d_{x^{2}-y^{2}}$ orbitals is near that of the optimal doped cuprates, $d$-wave HTSC with strongly enhanced $T_{c}$ will be achieved in this layer, and the top layer will also acquire SC below $T_{c}$ through proximity.
@@
Based on the above general consideration, we first conduct the following simplified model study including only the $3d_{x^{2}-y^{2}}$-orbital, with the $3d_{z^{2}}$ orbital only viewed as a source which tunes the total electron number. The following widely adopted single $3d_{x^{2}-y^{2}}$-orbital bilayer $t-J-J_{\perp}$ model [79, 80, 82, 92, 93, 95, 96, 100] is adopted,
@@
$H=$displaystyle-- $t_{\parallel}\sum_{\left\langle{}i,j\right\rangle,\mu,\sigma}\left(c^{{\dagger}}_{i\mu\sigma}c_{j\mu\sigma}+\mathrm{h.c.}\right)+\sum_{i,\mu}\epsilon_{\mu}n_{i\mu}$ (1)
@@
$+$ $J_{\parallel}\sum_{\left\langle{}i,j\right\rangle,\mu}\bm{S}_{i\mu}\cdot\bm{S}_{j\mu}+\tilde{J}_{\perp}\sum_{i}\bm{S}_{it}\cdot\bm{S}_{ib}.$
Here $c^{{\dagger}}_{i\mu\sigma}$ creates an electron at site $i$ in the layer $\mu$ (=top ($t$)/bottom ($b$)) with spin $\sigma$, and $n_{i\mu}$ or $\bm{S}_{i\mu}$ denote the corresponding electron number or spin operator. Only NN- bond $\langle i,j\rangle$ is considered in the summation. The $\epsilon_{\mu}$ is introduced to control the filling fractions of the two layers under $\bm{\varepsilon}$, with $\epsilon_{t}-\epsilon_{b}=\varepsilon$. However, as the total particle number of the $d_{x^{2}-y^{2}}$ electrons under given $\bm{\varepsilon}$ is unknown, we have to assume the ratio $r:1$ between the electron number flowing from the $3d_{z^{2}}$ orbitals and that flowing from $3d_{x^{2}-y^{2}}$ orbitals in the top layer when solving the model with the standard slave-boson mean-field (SBMF) theory [145]. Due to reason of DOS, we assume this ratio to be $2:1$, with details provided in the Supplementary Information (SI) [146]. Nevertheless, the concrete value of this ratio turns out not to obviously affect the results (see the SI [146]). The filling fractions are fixed under this assumption in the SBMF study. To capture the quantum fluctuation beyond mean-field, the density matrix renormalization group (DMRG) [147, 148] method is also employed, whose results are qualitatively consistent with those of the SBMF study. See details in the METHODS and SI [146].
We set $t_{\parallel}=1$ as the energy unit and $J_{\parallel}=0.4t_{\parallel}$ in our study. $\tilde{J}_{\perp}=(1-\delta_{tz})\times 1.3J_{\parallel}$ is applied in our SBMF study. The results are shown in Fig. 3. Fig. 3 (a) shows the amplitude and symmetry of the ground-state pairing gap as function of the bottom-layer $3d_{x^{2}-y^{2}}$ electron number $n_{bx}$, whose value enhances with $\bm{\varepsilon}$. It is shown that when $\bm{\varepsilon}$ or $n_{bx}$ enhances, the pairing amplitude $\tilde{\Delta}$ decays first and then increases. When $n_{bx}=0.5$, the ground state is interlayer $s$-wave SC. When $n_{bx}\geq 0.53$, the ground state is an intralayer $d$-wave SC, with the dominant pairing limited in the bottom layer. It is inspiring that with the enhancement of $n_{bx}$ in this regime, the $\tilde{\Delta}$ enhances promptly, similar to the case in the overdoped cuprates, wherein the enhancement of the filling fraction promptly enhances the pairing strength. The pairing configurations of the two different pairing symmetries are illustrated in Fig. 3 (c-d).
@@
On the above, we have adopted $\tilde{J}_{\perp}=\alpha J_{\perp}(1-\delta_{tz})$ with $\alpha=1$, where $\delta_{tz}$ denotes the hole density of the top-$3d_{z^{2}}$ orbital. For the reduced $\alpha$, only the low-$n_{bx}$ regime accommodating the interlayer $s$-wave pairing in Fig. 3 (a, b) shrinks but the high-$n_{bx}$ regime accommodating the intralayer $d$-wave SC is not affected because the intralayer pairing is blind to $\tilde{J}_{\perp}$. Furthermore, assuming different ratios between the changes of the filling fractions of the two top-layer $E_{g}$ orbitals turns out to yield similar results when expressed as functions of $n_{bx}$, as the dominant pairing under strong $\bm{\varepsilon}$ is the intra-bottom-layer pairing, which is blind to the filling fraction of the top layer. See the SI for details [146].

### v3

@@ Abstract
Recently, high-temperature superconductivity (HTSC) is found in the La3Ni2O7/SrLaAlO4 ultrathin film with critical temperature $T_{c}$ above the McMillan limit at ambient pressure (AP). It is eager to enhance $T_{c}$ of La3Ni2O7 at AP. We propose that a perpendicular electric field strongly enhances $T_{c}$ in the single-bilayer film of La3Ni2O7 at AP. Under electric field, the layer with lower potential energy will accept electrons flowing from the other layer to fill in the Ni-$3d_{x^{2}-y^{2}}$ orbitals, as the nearly half-filled Ni-$3d_{z^{2}}$ orbital cannot accommodate more electrons. With the enhancement of the filling fraction in the $3d_{x^{2}-y^{2}}$ orbitals in this layer, the interlayer $s$-wave pairing is suppressed, but the intralayer $d$-wave pairing in this layer is strongly enhanced. We numerically verify this idea and yield that an imposed voltage of about $0.1\sim 0.2$ volt between layers is enough to realize liquid-nitrogen-temperature HTSC in this single bilayer at AP. Our results appeal for experimental verification.
@@
The discovery of superconductivity (SC) with critical temperature $T_{c}$ above the boiling point of liquid nitrogen ($\approx 77\text{ K}$) in the pressurized La3Ni2O7 [Wang2023LNO, YuanHQ2023LNO, Wang2023LNOb, wang2023LNOpoly, wang2023la2prnio7, zhang2023pressure, zhou2023evidence, wang2024bulk, li2024pressure] has attracted great interests [Fukamachi2001, khasanov2024pressure, chen2024evidence, dan2024spin, chen2024electronic, Wang2022LNO, Kakoi2024, xie2024neutron, gupta2024anisotropic, ren2024resolving, feng2024unaltered, meng2024density, fan2024tunn, xu2024pressure, LI2024distinct, liu2024electronic, yashima2025microscopic, khasanov2025oxygen, yang2024orbital, wang2023structure, cui2023strain, Li2024ele, li2024distinguishing, zhou2024revealing, wang2025chemical, Chen2024poly, Dong2024vis, Li2024design, puphal2024unconven, zhu2024superconductivity, zhang2023superconductivity, huang2024signature, li2023trilayer, zhang2020intertwined, xu2024origin, du2024correlated, li2019superconductivity, lee2023linear, nomura2022superconductivity, gu2022superconductivity, sui2023rno, YaoDX2023, Dagotto2023, cao2023flat, zhang2023structural, huang2023impurity, geisler2023structural, rhodes2023structural, zhang2023la3ni2o6, yuan2023trilayer, li2024la3, geisler2024optical, li2017fermiology, wang2024non, chen2024tri, ZhangGM2023DMRG, Werner2023, shilenko2023correlated, WuWei2023charge, chen2023critical, ouyang2023hund, heier2023competing, wang2024electronic, botzel2024theory, WangQH2023, YangF2023, lechermann2023, Kuroki2023, HuJP2023, lu2023bilayertJ, oh2023type2, liao2023electron, qu2023bilayer, Yi_Feng2023, jiang2023high, zhang2023trends, qin2023high, tian2023correlation, jiang2023pressure, lu2023sc, kitamine2023, luo2023high, zhang2023strong, pan2023rno, sakakibara2023La4Ni3O10, lange2023mixedtj, yang2023strong, lange2023feshbach, kaneko2023pair, fan2023sc, wu2024deconfined, zhang2024prediction, zhang2024s, Yang2024effective, zhang2024electronic, yang2024decom, ryee2024quenched, Lu2024interplay, Ouyang2024absence, labollita2024, zhang2024emergent, Leonov2024Electronic, labollita2024assessing, ni2025spin, Yi2024nature, LaBollita2024electronic, jiang2024theory, chen2024non, zhang2024magnetic, lin2024magnetic, qin2024intertwined, Leonov2024Electronicc, liu2024growth, liu2024decoupling, tian2024spin, yin2025spmep, PhysRevB.111.104505, kaneko2025tj, Ji2025StrongCouplingLimit, Wang_2025, haque2025dft, shi2025theoretical, Ko2024signature, zhou2025ambient, liu2025superconductivity, yue2025correlated, li2025photoemission, bhatt2025resolving, shao2025band, shi2025effect, le2025landscape, Daoxin_Yao2025, wang2025electronic, huang2025spm, geisler2025electronic]. This discovery has sparked the exploration of high-temperature SC (HTSC) in Ruddlesden-Popper phase multilayer nickelates, resulting in the discovery of SC in the pressurized La4Ni3O10 [zhu2024superconductivity, zhang2023superconductivity, huang2024signature, li2023trilayer, zhang2020intertwined, xu2024origin, du2024correlated], which together with the previously synthesized infinite-layer nickelates Nd1-xSrxNiO2 [li2019superconductivity, lee2023linear, nomura2022superconductivity, gu2022superconductivity] have established a new family of SC other than cuprates and iron-based superconductors. However, the high pressure (HP) circumstance not only strongly hinders the experimental detection of the samples but also brings difficulties in the application of SC in industry. Very recently, the La3Ni2O7 ultrathin film with a few layers of unit cell grown on the SrLaAlO4 (SLAO) substrate has been grown by two different groups independently and SC with $T_{c}$ above the McMillan limit ($\approx$40 K) has been detected at ambient pressure (AP)[Ko2024signature, zhou2025ambient, liu2025superconductivity], allowing various experimental investigation of the pairing mechanism in this material, attracting a lot of interests [yue2025correlated, li2025photoemission, bhatt2025resolving, shao2025band, shi2025effect, le2025landscape, Daoxin_Yao2025, wang2025electronic, huang2025spm, geisler2025electronic]. It is now eager to enhance the $T_{c}$ of this material at AP. Here we propose a viable approach to realize $T_{c}$ above the boiling point of liquid nitrogen in the La3Ni2O7 single-bilayer film at AP.
@@
Presently, the pairing mechanism in the La3Ni2O7, either in the bulk material under HP [WangQH2023, YangF2023, lechermann2023, Kuroki2023, HuJP2023, lu2023bilayertJ, oh2023type2, liao2023electron, qu2023bilayer, Yi_Feng2023, jiang2023high, zhang2023trends, qin2023high, tian2023correlation, jiang2023pressure, lu2023sc, kitamine2023, luo2023high, zhang2023strong, pan2023rno, sakakibara2023La4Ni3O10, lange2023mixedtj, yang2023strong, lange2023feshbach, kaneko2023pair, fan2023sc, wu2024deconfined, zhang2024electronic, yang2024decom, ryee2024quenched, Lu2024interplay, Ouyang2024absence, yin2025spmep, PhysRevB.111.104505, kaneko2025tj, Ji2025StrongCouplingLimit] or in the ultrathin film at AP [yue2025correlated, shao2025band, le2025landscape, huang2025spm], is still under debate. Density-functional-theory (DFT) based first-principle calculations have suggested that the low-energy orbitals are mainly Ni-$3d_{z^{2}}$ and $3d_{x^{2}-y^{2}}$, which are nearly half- and quarter- filled [sui2023rno, YaoDX2023, Dagotto2023, cao2023flat, zhang2023structural, huang2023impurity, geisler2023structural, rhodes2023structural, zhang2023la3ni2o6, geisler2024optical]. Various experiments have revealed the strongly-correlated characteristic of the material [chen2024electronic, Kakoi2024, fan2024tunn, LI2024distinct, liu2024electronic, yang2024orbital, Li2024ele, li2024distinguishing]. Particularly, the optical study reveals significant reduction of the electron kinetic energy which places the system in the proximity of the Mott phase [liu2024electronic]; the angle-resolved photoemission spectroscopy uncovers strong band renormalization caused by electron correlation [yang2024orbital]; the linearly temperature-dependent resistivity suggests “strange-metal” behavior [YuanHQ2023LNO]. Therefore, we take a strong-coupling view of the system. Under the strong Hubbard repulsion, the nearly half-filled $3d_{z^{2}}$ electrons can almost be viewed as localized spins. Therefore, the main carrier of SC should be the $3d_{x^{2}-y^{2}}$ electrons, which subject to the in-plane superexchange interaction just mimic the 50% hole-doped cuprates. However, it is a problem how HTSC can emerge under such a high doping level. The key physics lies in the important role played by the $3d_{z^{2}}$ orbitals. Through strong interlayer perpendicular superexchange, the $3d_{z^{2}}$ electrons form interlayer pairing. The interlayer perpendicular superexchange or the interlayer pairing of the $3d_{z^{2}}$ electrons can be transmitted to the $3d_{x^{2}-y^{2}}$ electrons through the Hund’s rule [lu2023bilayertJ, oh2023type2, qu2023bilayer, tian2023correlation, zhang2023strong, pan2023rno, lange2023mixedtj, yang2023strong, lange2023feshbach, kaneko2023pair, wu2024deconfined, Lu2024interplay, Ouyang2024absence] or the nearest-neighbor (NN) hybridization [ZhangGM2023DMRG, Yi_Feng2023, qin2023high, tian2023correlation, luo2023high, kaneko2023pair, yang2024decom, Lu2024interplay, Ouyang2024absence] or both. Under such view, the role of pressure in enhancing the $T_{c}$ lies in the enhancement of the interlayer perpendicular superexchange, the inter-orbital hybridization, or both.
@@
Figure 1: Schematic diagrams of the model. (a) Schematic diagram for the dominant hopping integrals and superexchange interactions between the $E_{g}$ orbitals in La3Ni2O7. (b) Schematic diagram illustrating that the Hund’s rule coupling transmits the interlayer perpendicular superexchange interaction $J_{\perp}$ between the $3d_{z^{2}}$ orbitals to the effective one $\tilde{J}_{\perp}$ between the $3d_{x^{2}-y^{2}}$ orbitals.
@@
In this work, we propose an alternative approach to realize HTSC with $T_{c}$ above the boiling point of liquid nitrogen in the ultrathin film of La3Ni2O7 at AP. Here we consider the thinnest limit, i.e. a single bilayer film of La3Ni2O7, and realize the goal by introducing charge transfer with a perpendicular electric field, which let the electrons flow from the high-energy layer to the low-energy layer, similar to the mechanism for the spontaneous charge transfer in oxide heterostructures [PhysRevB.95.205131, PhysRevLett.123.117201, PhysRevB.98.155103, PhysRevB.108.054441, mondal2023modulation, doi:10.1021/acs.nanolett.4c02087]. The external electric field based approach avoids introducing disorder as in chemical doping [jiang2021electronic] or exhibiting orbital selectivity based on symmetry [sohn2024observation], demonstrating exceptional performance in the field of twisted multilayer graphene materials [cao2020tunable, zhang2024observation, zhou2025gate]. We can impose a perpendicular electric field, say pointing upward, in this single bilayer, so that electrons from the top layer will flow to the bottom layer. These electrons will fill the $3d_{x^{2}-y^{2}}$ orbitals in the bottom layer as the nearly half-filled $3d_{z^{2}}$ orbitals there cannot accommodate more electrons. The enhancement of the bottom-layer $3d_{x^{2}-y^{2}}$ electron number will first suppress the interlayer $s$-wave SC due to mismatch of the electron numbers between the two layers, and then promptly lead to the intra-bottom-layer $d$-wave SC with strongly enhanced $T_{c}$. To test this idea, we have performed a combined simplified single orbital study and a comprehensive two orbital study, which consistently yield that a voltage of experimentally achievable levels (around $0.1\sim 0.2$ volt predicted by the mean-field calculations) between the two layers is enough to induce $d$-wave SC with $T_{c}$ above the boiling point of liquid nitrogen in the bottom layer. Intriguingly, the $d$-wave SC carried by the bottom layer $3d_{x^{2}-y^{2}}$ electrons coexists with the interlayer $s$-wave pseudo gap carried by the $3d_{z^{2}}$ electrons in the mixing ratio of $1:\mathrm{i}$, breaking time-reversal symmetry. Our proposal potentially provides a viable approach to realize HTSC with $T_{c}$ above the boiling point of liquid nitrogen in the single bilayer film of La3Ni2O7.
@@ I.1 Consideration and a Simplified Study
The La3Ni2O7 ultrathin film grown on the SLAO substrate form a bilayer square lattice [bhatt2025resolving, yue2025correlated]. As illustrated in Fig. 1 (a), the leading hopping integrals are the interlayer hopping of the $3d_{z^{2}}$ electrons $t_{\perp}$ and the intralayer NN hopping of the $3d_{x^{2}-y^{2}}$ electrons $t_{\parallel}$. Under strong Hubbard $U$, these hopping terms can induce the effective superexchange $J_{\perp}$ and $J_{\parallel}$ through $J\approx\frac{4t^{2}}{U}$. Under the Hund’s rule coupling $J_{H}$, the spins of the two orbitals are inclined to be parallel aligned, as illustrated in Fig. 1 (b), which partly transmits the interlayer perpendicular superexchange $J_{\perp}$ between the $3d_{z^{2}}$ orbitals to the $3d_{x^{2}-y^{2}}$ orbitals as $\tilde{J}_{\perp}=\alpha J_{\perp}$ with $\alpha\in(0,1)$ describing the efficiency of this process and related to the strength of Hund’s coupling $J_{H}$. In addition, there exists intralayer NN- hybridization $t_{xz}$ between the two orbitals. As shown in Fig. 2 (a, b), the nearly quarter-filled $3d_{x^{2}-y^{2}}$ electrons subject to $J_{\parallel}$ and $\tilde{J}_{\perp}$ form interlayer-dominant pairing [lu2023bilayertJ].
[…]
@@
Figure 3: The SBMF results for the single-orbital model. (a) The pairing amplitude $\tilde{\Delta}$ (in unit of $t_{\parallel}$) as function of the bottom-layer particle number per site $n_{bx}$. Different pairing symmetries are distinguished by color. (b) The $T_{c}$ as function of $n_{bx}$, in comparison with $0.42\tilde{\Delta}$ for the $d$-wave and $0.9\tilde{\Delta}$ for the $s$-wave regime. Inset: the spinon pairing temperature $T_{\mathrm{pair}}$ and the holon condensation temperature $T_{\mathrm{BEC}}$ as function of $n_{bx}$. In (a,b), we set $J_{\parallel}=0.4t_{\parallel}$ and $\tilde{J}_{\perp}=(1-\delta_{tz})\times 1.3J_{\parallel}$. (c)-(d) The pairing configurations of the $s$-wave and $d$-wave, respectively.
[…]
@@
Based on the above general consideration, we first conduct the following simplified model study including only the $3d_{x^{2}-y^{2}}$-orbital, with the $3d_{z^{2}}$ orbital only viewed as a source which tunes the total electron number. The following widely adopted single $3d_{x^{2}-y^{2}}$-orbital bilayer $t-J-J_{\perp}$ model [lu2023bilayertJ, oh2023type2, qu2023bilayer, zhang2023strong, pan2023rno, lange2023mixedtj, yang2023strong, wu2024deconfined] is adopted,
@@
$H=$ - $t_{\parallel}\sum_{\left\langle{}i,j\right\rangle,\mu,\sigma}\hat{P}\left(c^{{\dagger}}_{i\mu\sigma}c_{j\mu\sigma}+\mathrm{h.c.}\right)\hat{P}+\sum_{i,\mu}\epsilon_{\mu}n_{i\mu}$ (1)
@@
$+$ $J_{\parallel}\sum_{\left\langle{}i,j\right\rangle,\mu}\mathbf{S}_{i\mu}\cdot\mathbf{S}_{j\mu}+\tilde{J}_{\perp}\sum_{i}\mathbf{S}_{it}\cdot\mathbf{S}_{ib}.$
[…]
@@
[…]